how many pounds of a 6 percent solution and water should be mixed together to obtain 150lbs of a 5 percent solution?
Answer by stanbon(75887)

how many pounds of a 6 percent solution and water should be mixed together to obtain 150lbs of a 5 percent solution?
------
Equation:
active + active = active
0.06x + 0*(150-x) = 0.05*150
---
Multiply thru by 100 to get:
6x + 0 = 5*150
---
x = 5*25
x = 125 lbs (amt. of 6% solution needed)
150-x = 25 lbs (amt of water needed)
==========================================
Cheers,
Stan H.